OpenStudy (anonymous):

completeing the square: x^2-10x-5=9

OpenStudy (anonymous):

You first get the constant numbers to the opposite side of the x^2, so get the -5 to the other side.

OpenStudy (anonymous):

x^2 - 10x = 14 do you understand this, so far?

OpenStudy (anonymous):

got it!!!!!! thank you so much!

OpenStudy (anonymous):

what did you get for an answer?

OpenStudy (anonymous):

x=+- square root of 39?

OpenStudy (anonymous):

x=5+_ square root of 39

OpenStudy (anonymous):

x^2 - 10x + 25 = 14 + 25 (x - 10)^2 = 39 x - 10 = +-sqrt(39) x = +- sqrt(39) + 10

OpenStudy (anonymous):

whered you get the 10 from?

OpenStudy (anonymous):

it factors to (x-5)^2

OpenStudy (anonymous):

OOPS you are correct....

OpenStudy (anonymous):

haha no worries

OpenStudy (anonymous):

but then you need to add 5 to both sides... x = +-sqrt(39) + 5 sorry I was texting and doing math at the same time.

OpenStudy (anonymous):

blexting can you help with a few more questions?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

M: 2/m + 3/m+9= 11/4

OpenStudy (anonymous):

You need to find a common denominator.. is the + 9 in the denominator or all by itself?

OpenStudy (anonymous):

m+9

OpenStudy (anonymous):

So the denominators are m and m + 9 and 4 so your common denominator is 4m(m+9) do you agree?

OpenStudy (anonymous):

yes i agreee!!!

OpenStudy (anonymous):

so we multiply everything by that correct?

OpenStudy (anonymous):

top and bottom?

OpenStudy (anonymous):

8(m+9)+12m =......

OpenStudy (anonymous):

Because you have denominators and an equal sign. you need to multiply each term by your common denominator. 4m(m+9)*2/m + 4m(m+9) * 3/(m+9) = 4m(m+9)*11/4

OpenStudy (anonymous):

Cancel the denominators into the numerators

OpenStudy (anonymous):

4(m+9)*2 + 4m*3 = m(m+9)*11

OpenStudy (anonymous):

8(m+9) + 12m = 11m(m+9)

OpenStudy (anonymous):

8m + 72 + 12m = 11m^2 + 99m

OpenStudy (anonymous):

Because you have an equal sign and an x^2 term you need to put the other side = to 0.

OpenStudy (anonymous):

0 = 11m^2 + 79m - 72

OpenStudy (anonymous):

I would use the quadratic formula here.. Do you agree with my steps so far?

OpenStudy (anonymous):

0 = (11m + 9)(m +8) m = -9/11 or -8 sorry it factors.

OpenStudy (anonymous):

im confuseddd ahhhhhhh i got 10.5

OpenStudy (anonymous):

i think i got it!

OpenStudy (anonymous):

omg got it!

OpenStudy (anonymous):

Great...

OpenStudy (anonymous):

i <3 you

OpenStudy (anonymous):

wanna be our math tutor for the night? i have another problem :D

OpenStudy (anonymous):

Sure... for awhile.